### v1

@@ Abstract
This paper examines the safety performance of the Waymo Driver, an SAE level 4 automated driving system (ADS) used in a rider-only (RO) ride-hailing application without a human driver, either in the vehicle or remotely. ADS crash data was derived from NHTSA’s Standing General Order (SGO) reporting over 7.14 million RO miles through the end of October 2023 in Phoenix, AZ, San Francisco, CA, and Los Angeles, CA. This study is one of the first to compare overall crashed vehicle rates using only RO data (as opposed to ADS testing with a human behind the wheel) to a human benchmark that also corrects for biases caused by underreporting and unequal reporting thresholds reported in the literature. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.78 IPMM for the human benchmark, an 85% reduction or a 6.8 times lower rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.85 IPMM for the human benchmark, a 57% reduction or 2.3 times lower rate. Police-reported and any-injury-reported crashed vehicle rate reductions for the ADS were statistically significant when compared in San Francisco and Phoenix as well as combined across all locations. The comparison in Los Angeles, which to date has low mileage and no reported events, was not statistically significant. In general, the Waymo ADS had a lower any property damage or injury rate than the human benchmarks. Given imprecision in the benchmark estimate and multiple potential sources of underreporting biasing the benchmarks, caution should be taken when interpreting the results of the any property damage or injury comparison. Together, these crash-rate results should be interpreted as a directional and continuous confidence growth indicator, together with other methodologies, in a safety case approach.
@@
1 Introduction
The safety of automated vehicles is a complex and evolving topic. The growing consensus, based on decades of experience in related safety-critical fields, is that a safety case approach is a suitable and structured method to demonstrate safety of an automated driving system (ADS), like the Waymo Driver (Favaro et al., 2023a). The safety case approach described in Favaro et al. (2023a) describes three complementary views on safety determination: a layered approach, a credible approach, and a dynamic approach to safety. The layered approach comprises a goal of demonstrating absence of unreasonable risk, where risks are decomposed into architectural, behavioral, and operational layers and appropriate acceptance criteria are set. The credible approach describes a case credibility assessment (CCA), where the credibility of the argument and evidence in the safety case can be evaluated. Finally, the dynamic approach describes the safety determination lifecycle where safety is assessed as an emergent property, predictions, and confidence growth over time. The safety methodologies described in Webb et al. (2020) can be seen as a “collective set of evidence used to populate the safety case” (Favaro et al., 2023a).
@@ 1 Introduction
This paper describes crash results from an in-field deployment of a Rider-Only (RO) ADS ride-hailing service as part of the safety determination lifecycle. The in-field crash results may serve as one important factor in confirming design elements and predictions done in earlier iterations of this safety determination lifecycle, as shown in Figure 1 (Favaro et al., 2023a). While a system is being developed and before starting RO operations or when considering an update to an existing RO deployment, only prospective methods which predict expected performance can be applied. In a prospective method, for example in simulated deployments (Webb et al., 2020) or as described in Favaro et al. (2023b), simulation is used to predict an ADS crash rate and compare that rate to a benchmark. In this approach, benchmarks can be set for different severity levels that are determined based on a crash severity model, as outcomes such as injury are not available from simulation. Retrospective analyses based on crash outcomes like presented in this paper can be used to complement and/or confirm these prospective methods that are used in an ADS readiness determination (Webb et al., 2020).
@@
Figure 1: Visualization of Waymo’s Safety Determination Life Cycle (adapted from Figure 6, Favaro et al. (2023a))
@@
Retrospective examination of crash performance against a human benchmark is one of many safety determination activities that comprise an ADS safety case. It is well understood, however, that the human driving population engages in behaviors that increase risk of crashes, such as driving while intoxicated, distracted, or drowsy. Therefore, an ADS with a crash involvement rate lower than the general driving population does not by itself allow for a determination that the ADS has an absence of unreasonable risk. Assessment methodologies that use other benchmarks are important in addition to retrospective crash involvement. One example is the Collision Avoidance Testing (CAT) program that uses virtual simulation to compare the ADS crash performance in nominal (i.e., not degraded) responder role conditions to a Non-Impaired Eyes On Conflict (NIEON) (Kusano et al., 2022). The NIEON model is fit to human near-crash data where the driver had their eyes on the conflict and is not an achievable state for real human drivers, who will always have some level of not looking at the road. The broader safety case approach, with its goal to demonstrate an absence of unreasonable risk, considers a broad range of potential issues that can come from the ADS ride-hailing operations, including measures of the ADS’s drivership, the realization of expected driving behaviors that position the ADS as a good citizen of the road. Evaluation of drivership can extend beyond crash risk.
For over 50 years, the study of motor vehicle safety has relied on crash reporting databases to understand trends and develop countermeasures to improve traffic safety. Examples of these databases in the United States are the Fatal Analysis Reporting System (FARS), the Crash Report Sampling Systems (CRSS), and the Crash Investigation Sampling System (CISS). Each of these data sources are maintained by the National Highway Traffic Safety Administration (NHTSA). Estimates of Vehicle Miles Traveled (VMT) are also maintained by federal and local transportation agencies, such as the Federal Highway Administration’s “Highway Statistics” annual reporting (Federal Highway Administration, 2023). Putting these types of crash reporting and VMT estimates together, public agencies such as NHTSA often examine the trends in the rate of different severities of crashes. For example, in 2021 (the most recent year that complete data is available and analyzed), the rate of fatalities per 100 million VMT was 1.37 and the rate of injuries per 100 million VMT was 80 (Stewart, 2023). Analyzing the trends in crash rates over time and in different conditions (e.g., urban vs rural areas) is possible because there is a consistent reporting standard that changes little from year-to-year.
@@
ADS manufacturers have also been required to report crashes for some time, albeit with a different reporting standard than in police-reported crash databases. The California Department of Motor Vehicles was one of the first agencies to collect and publish data on crashes involving ADS-equipped vehicles, in their case those operating in California. More recently, the NHTSA Standing General Order (SGO) 2021-01 has required ADS developers to report all crashes meeting certain criteria, which are then published publicly since July 2021 (National Highway Traffic Safety Administration, 2023). Because ADS are emerging technologies and are currently deployed in small numbers compared to human-operated vehicles, the reporting requirements for ADS manufacturers is to report any crashes that the manufacturer knows to have occurred or that are alleged and that result in “any property damage, injury, or fatality”, occur on a public road, and in which the ADS was engaged at any time during the 30 seconds immediately prior to the crash through the conclusion of the crash (National Highway Traffic Safety Administration, 2023).
The observed crash and injury reduction following the introduction of ADS technology is referred to as retrospective safety impact. A valid retrospective safety impact assessment requires accurate calibration of benchmark human and ADS crash and VMT data sources to make a valid comparison. Different approaches are needed depending on the nature of the human benchmark and ADS data sources. For example, in a reported crash database, such as data derived from police reports, there is a need to apply some underreporting adjustment [^1] For ease of writing, we will refer to “underreporting” in reported crash data to include both crashes that meet the reporting threshold but were not reported and crashes that do not meet the reporting threshold. See the discussion in Scanlon et al. (2023) for a more thorough discusion of these topics, and how the Blincoe et al. (2022) and Blanco et al. (2016) underreporting adjustments differed in regards to these two dimensions. if the ADS data includes any property damage. This underreporting adjustment to police-reported crashes is the approach Scanlon et al. (2023) took to estimate crash rates for any property damage. Another human data source used for generating a benchmark is Naturalistic Driving Study (NDS) databases, which equip vehicles with sensors such as accelerometers and cameras to record driving. Using a combination of algorithmic triggers (e.g., hard braking or swerving) plus manual review to find true positive crashes based on the video data, an observed crash rate can be found. NDS generally have less advanced sensors than those that are equipped on a SAE level 4 ADS vehicle, which may induce, for example, blind spots in the camera or sensor views. Other NDS data sources may have even less data recorded, such as those from dash cameras that may only have a front camera view. The less advanced sensors also necessitate the approach of using kinematic triggers followed by manual review of the video data. This approach may lead to missing property damage crashes that do not meet the kinematic triggers (e.g., low speed contacts while the ego vehicle is stationary). To adjust for potential missed crashes, Flannagan et al. (2023) used a mixture of insurance, telematics, and NDS data to estimate any property damage crash rate.
@@
Given public accessibility of both human and ADS crash and VMT data, researchers have the opportunity to begin evaluating ADS’s safety performance. Working with these public crash and VMT data sources, however, has certain challenges. Several reviews of the literature have found that many of the early studies comparing ADS and human benchmarks introduced biases (Young, 2021; Scanlon et al., 2023). Scanlon et al. (2023) describes errors and biases that should be accounted for when attempting to compare human and ADS crash rates: (1) information biases (surveillance bias caused by reporting thresholds, reporting bias caused by underreporting), and (2) biases from analytical choices (selection bias caused by lack of operational design domain, ODD, matching). A large number of human crashes are not reported to police or are reported but no police report is filed (reporting bias). NHTSA estimates that 60% of property-damage-only and 32% of injury crashes are not reported to police (Blincoe et al., 2022). The reporting requirement of “any property damage”, without any lower limit for property damage for ADS crash data, is a lower threshold than human crash data sources such as CRSS or state police report data that require a police report to be filed in order to appear in these data sources (surveillance bias). In contrast, ADS fleet operators are required to report almost every crash and the large number of sensors combined with operational procedures result in almost all contacts being discovered and reported. Finally, the driving environment and/or vehicle characteristics of a datasource can impact the crash rate (selection bias). Factors such as population density (rural vs. urban), roadway type (surface street vs freeway), and vehicle use type (commercial heavy vehicles, passenger vehicles, ride hail vehicle) can all affect crash rates.
@@
Generating a valid comparison of human and ADS crash rates that adjusts for the above mentioned biases is of great importance at this point in time. Larger ADS deployments in recent years have led to inquiries from the general public, public officials, and researchers about the safety of current ADS fleets. Improving traffic safety is a goal held by many ADS stakeholders. All stakeholders have an interest in understanding the current state of ADS safety performance through as many objective measures as are available, which include retrospective crash rates. If biases are present in the data that either inflate or deflate the ADS or benchmark crash rates, the conclusions may lead to incorrect public perceptions of ADS safety and questionable policy recommendations or decisions. As pointed out in Scanlon et al. (2023, Table 1) and Young (2021), past studies have varying degrees of biases in their analyses. In addition, non-academic studies, either in the popular press or done by government agencies, are being done. All data sources, both from human and ADS sources, have limitations in the data available. Therefore, generating sets of crashes that are comparable between human benchmarks and ADS sources is complex and requires judgment based on recognized best practices. The analysis in this paper serves to present an analysis that uses publicly available ADS and benchmark data in order to advance the understanding of ADS safety so that stakeholders can make informed decisions. This analysis is also a step toward further harmonization and standardization of retrospective safety impact analysis for ADS fleets.
@@
The purpose of this paper is to present Waymo ADS crash rates from data reported according to NHTSA’s Standing General Order (SGO) and compare to benchmark rates that are as accurate as possible as derived from comparable human benchmark sources published in the literature.
The NHTSA SGO requires reporting of all crashes involving or alleged to have involved the ADS vehicle resulting in any amount of property damage, injury, or fatalities if the crashes occur on public roads and the ADS was engaged at any time during the 30 seconds immediately prior to the crash through the conclusion of the crash (National Highway Traffic Safety Administration, 2023). The SGO reporting is required in all states, and thus is the most complete crash reporting required for an ADS fleet operator. Comparisons were made for crashes involving any property damage or injury, police-reported, and any-injury-reported outcomes. As the SGO data is publically available, the analyses provided herein can be readily reproduced and developed by other researchers.
2 Methods
2.1 Selection of Waymo ADS Incidents
@@
The Waymo RO ride-hailing service in Phoenix, San Francisco, and Los Angeles operates in fixed geographic areas in these cities operating 24 hours a day and 7 days a week. The ODD includes non-limited access roads with speed limits up to 50 mph and parking lots without restrictions on maneuvers. The ODD does not include severe weather conditions, such as thick fog, heavy rain, or blowing sand but does include light rain or light fog. The ODD for the Waymo RO service has changed slightly over the years as the service expanded. The ODD features have been relatively unchanged in the past year, which also corresponds to the majority of the RO miles accumulated to date. Therefore, the Waymo RO crashes described in this section were compared to crashes that occurred in the same geographical areas. The selection of human benchmarks is discussed in detail in the following section.
As NHTSA SGO ADS crashes are a mandated and publicly available data source, Waymo’s NHTSA SGO reportable crashes became the basis for the incidents examined in this study [^2] Data is available for download at https://www.nhtsa.gov/laws-regulations/standing-general-order-crash-reporting. The NHTSA SGO requires reporting crashes where the ADS was engaged at any time (not throughout) the 30-second pre-crash and crash period, regardless of whether a human driver operated the vehicle during that period or whether the ADS was operating the vehicle in RO mode during some or all of that period. For this analysis, only Waymo SGO-reported crashes where the ADS was operating in RO mode were included for analysis. In the SGO data, collisions reported by Waymo were identified using the “Reporting Entity” field having the value “Waymo LLC”. The field “Driver / Operator Type” was set to “None” for RO operations, and set to “In-Vehicle (Commercial / Test)” when an autonomous specialist is monitoring the ADS behind the steering wheel. In those ADS crashes with an autonomous specialist present, it is difficult to disambiguate the contribution of the ADS driver and the autonomous specialist. Therefore, this analysis excluded those crashes with an autonomous specialist present behind the wheel, as RO crashes are most representative of the intended autonomous ride-hailing service.
Not-in-transport vehicles are excluded in the crashed vehicle count from the police-reported database human driver benchmarks used in this study. Any SGO-reported crash where the Waymo ADS vehicle was a vehicle not-in-transport were identified to better compare the ADS crashes to police crash report benchmarks. An “in-transport” vehicle is one that is traveling (moving or stopped) in the roadway. “Not-in-transport” vehicles are those that are parked in a designated parking spot out of the roadway. Therefore, including ADS vehicle crashes where the ADS vehicle is a not in-transport vehicle would inflate the ADS crash count relative to the benchmark. The ADS vehicle was considered not in-transport if the ADS vehicle was in the “park” gear and the ADS vehicle was in a designated parking area. This not in-transport status can happen when the ADS vehicle is parked awaiting its next trip instructions. Because the ADS software is running and in control of the vehicle, contacts while the ADS vehicle is not in-transport are reported as part of the NHTSA SGO reporting requirements. If the parking spot was on-street parking parallel to the traffic flow, the ADS vehicle must have been within 18 inches from the curb of the road as measured by the measured sensor data and curb location in the ADS map to be considered not-in-transport. Instances where the ADS vehicle was in “park” gear, but not near the curb were included as in-transport vehicles. The alternative, to include “not-in-transport” vehicle ADS crashes, would require also including “not-in-transport” vehicles in the benchmark, which are included in separate counts from in-transport vehicles in most police report databases, if at all. Additionally, it is unclear how a rate of crashed vehicles per VMT should be interpreted for not-in-transport vehicles, as parked vehicle crash risk is likely more related to amount of time parked rather than VMT. Underreporting for not-in-transport vehicles may also differ from in-transport vehicles, which was not accounted for the underreporting adjustments used in the benchmarks (Scanlon et al., 2023; Blincoe et al., 2022).
Additionally, the NHTSA SGO requires ADS fleet operators to report crashes where the ADS vehicle contributed or is alleged to have contributed (by steering, braking, acceleration, or other operational performance) to another vehicle’s physical impact with another road user or property involved in a reportable crash even if the ADS vehicle was not involved in the physical impact. One NHTSA SGO-reported crash (30270-6542) involved two human-driven vehicles that made contact while following behind a Waymo ADS vehicle. The Waymo vehicle was not contacted or damaged. Because the purpose of this paper is to compare crashed vehicle rates between ADS and human benchmarks, we excluded this crash from the NHTSA SGO in-transport vehicle category and subsequent categories. This single event was included in the SGO-reported count. The reason to exclude this type of crash where the ADS vehicle does not impact another vehicle is because vehicles that are not involved in an impact are not included in the crashed vehicle counts in the benchmark crash sources discussed in the following section. Therefore, including SGO-reported events where the ADS did not impact another vehicle or object would inflate the ADS rate compared to the benchmark.
In addition to examining all NHTSA SGO-reported crashes, results were also reported for a dataset that excluded minor crashes, which we will refer to as “SGO Crash Excluding Low Delta-V.” Although the benchmark data used in this study and described in detail in the next section was either adjusted for underreporting or used NDS data designed to find most contacts, there is still uncertainty in the benchmark data about the lower reporting threshold for a crash. This result of excluding low delta-V ADS crashes was presented alongside the larger all SGO-reported crashes group to be able to investigate the sensitivity of lower reporting threshold on the comparison to the benchmark. We defined these minor contacts as those where a crash reconstruction model found that both vehicles had a change in velocity (i.e., a delta-V) of less than 1.0 mph. The collision model used was an impulse-momentum model using the inertial properties of the ADS equipped vehicle and estimated inertial properties of the crash partner, as described in Scanlon et al. (2021). Each of the 7 crashes with fixed or non-fixed objects was examined individually to estimate a delta-V, discussed in more detail in the appendix. Of the 7 crashes with fixed or non-fixed objects, 5 were excluded for having a low delta-V. Crashes where the ADS was not-in-transport were also excluded from this group. The single crash involving a cyclist (30270-5456) was included in this group.
Next, police-reported and alleged-injury crashes were identified from the SGO-reported crashes. Police-reported crashes were identified using the “Law Enforcement Investigating?” field in the SGO report data. Any SGO-reported crashes that had “Yes” was considered police-reported. For any SGO-reported crashes with “Unknown” in the “Law Enforcement Investigating?” field, Waymo used internal records (e.g., field incident response communication, insurance claims information) to determine if a police report was filed. To determine any injury crashes, the “Highest Injury Severity Alleged” SGO field was used. Any crash that did not have “No Injuries Reported” was considered an any-injury-reported crash. In this dataset, there were 3 crashes with “Minor” injuries reported. Crashes where the ADS was not-in-transport were also excluded from police-reported and alleged any-injury-reported crashes.
A list of all crash incidents included for analysis in this paper is included in the Appendix. Each NHTSA SGO-reported RO crash is listed along with whether the crash was included in the in-transport vehicle, low delta-V, police-reported, or any-injury-reported outcome groups. Detailed narratives or other data is not included in this paper, as this information can be obtained from the NHTSA SGO reporting website.
2.2 Human Benchmarks
Table 1 lists the human benchmarks used to compare to the ADS crashes in this study. The ADS data is derived from the NHTSA SGO-reported crash data. Human crashed vehicle rates derived from all police-reported crashes with two different underreporting adjustments were reported in Scanlon et al. (2023). The Blincoe-adjusted estimate was derived from a combination of a national phone survey and comparing insurance and crash record data as reported in Blincoe et al. (2022). The Blanco-adjusted estimate used the proportion of unreported to police-reported crashes from the Strategic Highway Research Program 2 (SHRP-2) NDS reported in Blanco et al. (2016). Two other property damage benchmarks were considered. First, a benchmark derived from NDS from a population of ride-hailing drivers in San Francisco was described by Flannagan et al. (2023). Second, the overall crashed vehicle rate (excluding tire strikes) from the SHRP-2 NDS was reported in Blanco et al. (2016). The SHRP-2 NDS was conducted in multiple urban and rural areas of the country, so it is most appropriate to compare this benchmark to the aggregate ADS crashes. The two NDS-derived benchmarks both had a higher rate (by a factor of between 1.7 and 8 times) than the Blincoe-adjusted crashed vehicle rates reported in Scanlon et al. (2023). This large discrepancy between the Blincoe-adjusted and NDS rates suggests the Blincoe et al. (2022) property damage adjusted may not represent an any property damage reporting threshold. The Blanco et al. (2016) adjustment for property damage only is derived from NDS data, which as discussed above, is based on recorded events and not self-reported survey data as was used in the Blincoe et al. (2022) adjustment. We therefore used the Blanco et al. (2016) adjustment estimate as the primary comparison to the ADS crashes in this study. The human crash benchmarks reported in Scanlon et al. (2023) in Phoenix, San Francisco, and Los Angeles were combined proportional to the miles driven in the Waymo RO service and are shown in Table 2 as the “mileage blended average.” This mileage blended average was compared to the crashed vehicle rate aggregated in the three locations.
As noted in Scanlon et al. (2023), the crashed vehicle count from the police reported data only included in-transport vehicles. Therefore, the “all property damage” crashes from Scanlon et al. (2023) were compared with all SGO-reported ADS crashes, excluding those where the ADS vehicle was not-in-transport as described in the previous section. Because the NDS data sources described in Flannagan et al. (2023) and Blanco et al. (2016) did not specify if not in-transport NDS vehicles were excluded, we compared these NDS data to all reported SGO ADS crashes.
The second comparison made in this study was between ADS NHTSA SGO crashes with a police report filed and the observed human police-reported rate. Third, ADS NHTSA SGO crashes where any injuries were reported were compared to the human injury benchmark, adjusted for underreporting. Because police-reported and any-injury-reported crashes are relatively rare, most human NDS-based benchmarks have few of these crashes and do not attempt to estimate rates. Police-reported crash databases, on the other hand, aggregate large numbers of crashes over the entire driving fleet. Therefore, the only comparable benchmark data sources available were the police reported and any injury reported (adjusted for underreporting) rate estimates presented in Scanlon et al. (2023).
Table 1 shows the most valid comparisons between the ADS and human benchmark data sources. Table 2 lists all the human benchmark crashed vehicle rates and the underlying VMT the rates are based on.
@@ 2.2 Human Benchmarks
Any Property Damage or Injury
All NHTSA SGO Crash - In-Transport Vehicles
Scanlon et al. (2023) Any Property Damage or Injury Blanco Adjusted
All NHTSA SGO Crashes in San Francisco
Flannagan et al. (2023)
All NHTSA SGO Crash in all Locations
Blanco et al. (2016)
Police-Reported
All NHTSA SGO Crash - In-Transport Vehicles with Police Report
Scanlon et al. (2023) Police-reported Unadjusted
Any-injury-reported
All NHTSA SGO Crash - In-Transport Vehicles with Reported Injury
Scanlon et al. (2023) Any Injury Blincoe-Adjusted

### v2

@@ Abstract
Objectives: This paper examines the safety performance of the Waymo Driver, an SAE level 4 automated driving system (ADS) used in a rider-only (RO) ride-hailing application without a human driver, either in the vehicle or remotely.
Methods: ADS crash data were derived from NHTSA’s Standing General Order (SGO) reporting over 7.14 million RO miles through the end of October 2023 in Phoenix, AZ, San Francisco, CA, and Los Angeles, CA and compared to human benchmarks from the literature.
Results: When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Police-reported and any-injury-reported crashed vehicle rate reductions for the ADS were statistically significant when compared in San Francisco and Phoenix, as well as combined across all locations. The any property damage or injury comparison had statistically significant decrease in 3 comparisons, but also non-significant results in 3 other benchmarks. When excluding ADS crashes with a delta-V less than 1 mph (a measure of sensitivity to lower reporting threshold), about half of the ADS collisions were excluded resulting in comparisons that showed a large statistically significant reduction in all comparisons except for one comparison from San Francisco.
Conclusions: The statistically significant reductions in police-reported and any-injury-reported crash rates indicate a promising positive safety impact of ADS. The direction and significance of comparisons in the any property damage or injury outcome group are inconclusive due to difficulties in estimating a matching human benchmark. More research is needed on defining any property damage or injury benchmarks with clear lower reporting thresholds to reduce the systematic uncertainty in the benchmark rates. Together, these crash-rate results contribute to the continuous confidence growth, together with other methodologies, in a safety case approach.
@@
INTRODUCTION
@@
This paper describes crash results from an in-field deployment of a Rider-Only (RO) ADS ride-hailing service as part of the safety determination lifecycle. The in-field crash results may serve as one important factor in confirming design elements and predictions done in earlier iterations of this ADS safety determination lifecycle, as shown in Figure 1 (Favarò et al., 2023a). While a system is being developed and before starting RO operations or when considering an update to an existing RO deployment, only prospective methods which predict expected performance can be applied. In a prospective method, for example in simulated deployments (Webb et al., 2020) or as described in Favarò et al. (2023b), simulation is used to predict an ADS crash rate and compare that rate to a benchmark. In this approach, benchmarks can be set for different severity levels that are determined based on a crash severity model, as outcomes such as injury are not available from simulation. Retrospective analyses based on crash outcomes like presented in this paper can be used to complement and/or confirm these prospective methods that are used in an ADS readiness determination (Webb et al., 2020).
@@
Figure 1: Visualization of Waymo’s Safety Determination Life Cycle (adapted from Figure 6, Favarò et al. (2023a))
@@
Retrospective examination of crash performance against a human benchmark contributes to the continuous confidence growth portion of an ADS safety case. Retrospective studies cannot be used to make safety determinations prior to deployment. Therefore, an ADS with a crash involvement rate lower than the general driving population does not, by itself, allow for a determination that the ADS has an absence of unreasonable risk. Assessment methodologies that use other benchmarks are important in addition to retrospective crash involvement (Webb et al., 2020). The broader safety case approach considers a range of potential issues that can arise during ADS ride-hailing operations, including measures of the ADS’s drivership, the realization of expected driving behaviors that position the ADS as a good citizen of the road. Evaluation of drivership can extend beyond crash risk.
@@
The data sources most used to understand traffic safety historically have been derived from police-reported crash databases, like those maintained by the National Highway Traffic Safety Administration (NHTSA) and states. Recently, ADS manufacturers have been required to report crashes, albeit with a different reporting standard than that used in police-reported crash databases. Since July 2021, the NHTSA Standing General Order (SGO) 2021-01 has required ADS developers to report all crashes meeting certain criteria, which are then published publicly (National Highway Traffic Safety Administration, 2023). The reporting requirements for ADS manufacturers is to report any crashes that the manufacturer knows to have occurred or that are alleged to have occurred and that result in “any property damage, injury, or fatality”, occur on a public road, and in which the ADS was engaged at any time during the 30 seconds immediately prior to the crash through the conclusion of the crash (National Highway Traffic Safety Administration, 2023).
A retrospective safety impact assessment requires calibration of benchmark human and ADS crash and vehicle miles traveled (VMT) data sources to make a valid comparison. Given public accessibility of both human and ADS crash and VMT data, researchers have the opportunity to begin evaluating ADS safety performance. Working with these public crash and VMT data sources, however, has certain challenges. Several reviews of the literature have found that many of the early studies comparing ADS and human benchmarks introduced biases (Young, 2021; Scanlon et al., 2023).
One of the most important considerations when matching benchmark and ADS crash and VMT data sources is the inclusion criteria of the data sources being compared (i.e., to correct for surveillance bias). Figure 2 shows the conceptual relationship between different crash severity reporting thresholds relative to their frequency of occurrence. The distribution shows hypothetical frequency by crash severity (e.g., delta-V or impact speed). Below the frequency distribution, there are ranges of severity that correspond to different reporting thresholds for crash data sources. The NHTSA SGO data is of the “any property damage” reporting threshold. Most human crash databases such as NHTSA’s Crash Report Sampling Systems, CRSS, database and state crash databases are derived from crashes reported to police and/or through insurance claims. A subset of police-reported crashes can result in a reported injury (i.e., minor to fatal injuries), and a smaller subset of those any injury reported crashes result in fatal injuries. Given these different reporting thresholds, it is not appropriate to compare ADS crash data like the NHTSA SGO, which includes outcomes at the any property damage level, directly to human police-reported crash databases like national or state crash databases.
Figure 2: Conceptual Relationship Between Different Crash Severity Reporting Thresholds. The arrows below the graph show the hypothetical range of severities included in each outcome group.
@@
In addition to reporting thresholds, another form of bias comes from reporting bias due to underreporting. Many human crashes are not reported to police or are reported but no police report is filed. In the U.S.  60% of property-damage-only and 32% of injury crashes are not reported to police (Blincoe et al., 2023). In contrast, ADS fleet operators are required to report almost every crash and the large number of sensors combined with operational procedures result in almost all contacts being discovered and reported. Finally, the driving environment and/or vehicle characteristics of a data source can impact the crash rate (selection bias). Factors such as population density (rural vs. urban), roadway type (surface street vs freeway), and vehicle use type (commercial heavy vehicles, passenger vehicles, ride hail vehicle) can all affect crash rates.
Due to these potential challenges that introduce bias into comparisons of benchmark and ADS crash data, different approaches are needed to form valid comparisons depending on the nature of the human benchmark and ADS data sources. For example, if a benchmark that is derived from crashes reported to police is compared to ADS data that includes any property damage, there is a need to either apply some underreporting adjustment or restrict the ADS data to only include police-reported crashes. For ease of writing, we will use “underreporting” to refer to both crashes that meet the reporting threshold but were not reported and crashes that do not meet the reporting threshold. See the discussion in Scanlon et al. (2023) for a more thorough discussion of these topics. An underreporting adjustment to police-reported crashes is the approach Scanlon et al. (2023) took to estimate crash rates for any property damage. Alternatively, another human data source used for generating a benchmark is Naturalistic Driving Study (NDS) databases, which equip vehicles with sensors such as accelerometers and cameras to record driving. Using a combination of algorithmic triggers (e.g., hard braking or swerving) plus manual review to find crashes based on the video data, an observed crash rate can be found, such as reported in Blanco et al. (2016) and Flannagan et al. (2023).
@@
Generating a valid comparison of human and ADS crash rates that adjusts for the above mentioned biases is of great importance at this point in time. Larger ADS deployments in recent years have led to inquiries from the general public, public officials, and researchers about the safety of current ADS fleets. All stakeholders have an interest in understanding the current state of ADS safety performance through as many objective measures as are available, which include retrospective crash rates. If biases are present in the data that either inflate or deflate the ADS or benchmark crash rates, the conclusions may lead to incorrect public perceptions of ADS safety and questionable policy recommendations or decisions. As pointed out by Scanlon et al. (2023, Appendix A.1.2) and Young (2021), past studies have varying degrees of biases in their analyses. In addition, non-academic studies, either in the popular press or by government agencies, are being done. All data sources, both from human and ADS sources, have limitations. This paper aims to present an analysis that uses publicly available ADS and benchmark data in order to advance the understanding of ADS safety so that stakeholders can make informed decisions. This analysis is also a step toward further harmonization and standardization of retrospective safety impact analysis for ADS fleets.
In the field of traffic safety, the prevention of severe crashes is generally prioritized. One of the paradoxes of the study of crashes, however, is that severe outcomes are less frequent than low severity outcomes. This is true for several types of outcomes. For example, crashes with low dollar values of property damage outnumber high dollar values and minor injuries outnumber serious injuries. Given that ADS vehicles drive fewer miles relative to human-driven vehicles, the ADS crash data are overwhelmingly property damage only crashes. As discussed above, it is difficult to construct comparable benchmarks for any property damage crashes due to differences in reporting thresholds and underreporting. Because of the great public interest, this paper will attempt to construct a benchmark and compare it to ADS crashes for an any property damage or injury outcome despite the challenges. Because the most-studied human crash data is either at the police-reported or any-injury-reported outcome level (or greater), these data sources are more well understood and are likely to have less systematic biases present compared to any property damage or injury benchmarks. The drawback of police-reported and any-injury-reported outcome levels is that there will be fewer ADS crashes observed, leading to potentially larger confidence intervals. Because more serious outcome crashes often have different characteristics than less severe outcome crashes, it is possible that the safety impact and/or the confidence in the impact of an ADS may be different for different outcome levels. See the Appendix A.4 for further discussion on considerations for safety impact assessment for serious injury and fatal crashes.
@@
The purpose of this paper is to present Waymo ADS crash rates from data reported according to NHTSA’s SGO and compare to benchmark rates from comparable human benchmark sources published in the literature. Comparisons were made for crashes involving any property damage or injury, police-reported, and any-injury-reported outcomes. As the SGO data is publicly available, the analyses provided herein can be readily reproduced and developed by other researchers.
METHODS
Matching of Benchmark and ADS Data
Due to the data collection differences between human and ADS data, there is no set of data that represent the perfect comparison between human and ADS crash rates. Therefore, this paper aimed to develop several comparable sets of human and ADS crash rates using data published from the literature and NHTSA SGO data.
@@ METHODS
The Waymo RO ride-hailing service in Phoenix, San Francisco, and Los Angeles operates in fixed geographic areas in these cities operating 24 hours a day and 7 days a week. The ODD includes non-limited access roads with speed limits up to 50 mph and parking lots without restrictions on maneuvers. The ODD does not include severe weather conditions, such as thick fog, heavy rain, or blowing sand, but does include light rain or light fog. The ODD for the Waymo RO service has changed slightly over the years as the service expanded. The ODD features have remained largely consistent since 2022, which also corresponds to the majority of the RO miles accumulated during the study period (through October 2023). Therefore, the Waymo RO crashes described in this section were compared to crashes that occurred in the same geographical areas. The selection of human benchmarks is discussed in detail in the following section.
As NHTSA SGO ADS crashes are a mandated and publicly available data source, Waymo’s NHTSA SGO reportable crashes became the basis for the incidents examined in this study. The SGO reporting is required in all states, and thus is the most complete crash reporting required for an ADS fleet operator. Manufacturers do not report miles driven as part of the SGO data. Therefore, this study analyzed only Waymo RO crashes as these miles were the only ones available to the authors. All crashes involving Waymo operating in RO through the end of October 2023 were included. SGO data was downloaded from NHTSA’s website in December 2023.
Table 1 lists the ADS data and human benchmarks compared in this study and Table 2 lists the benchmark crash rates. Comparisons were grouped into three outcome groups: Any Property Damage or Injury, Police-reported, and Any-Injury-Reported outcomes. Details of each outcome group are discussed below.
@@
Any Property Damage or Injury All NHTSA SGO Crash - In-Transport Vehicles Scanlon et al. (2023) Any Property Damage or Injury Blincoe-Adjusted
NHTSA SGO Crashes - In-Transport Vehicles Excluding Low Delta-V (< 1 mph) Scanlon et al. (2023) Any Property Damage or Injury Blincoe-Adjusted
All NHTSA SGO Crashes in San Francisco Flannagan et al. (2023)
All NHTSA SGO Crash in all Locations Blanco et al. (2016)
Police-Reported All NHTSA SGO Crash - In-Transport Vehicles with Police Report Scanlon et al. (2023) Police-Reported Unadjusted
Any-Injury-Reported All NHTSA SGO Crash - In-Transport Vehicles with Reported Injury Scanlon et al. (2023) Any-Injury-Reported Blincoe-Adjusted
All NHTSA SGO Crash - In-Transport Vehicles with Reported Injury Scanlon et al. (2023) Any-Injury-Reported Observed